OpenStudy (anonymous):

Which fraction is equivalent to 75%

OpenStudy (anonymous):

3/4

OpenStudy (ajanijones):

Yep.

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

3/4 aw man.

OpenStudy (ajanijones):

Like saying 4 quarters = 4/4 =1 3/4 = 75% 2/4 = 50% 1/4 = 25%